Use algebraic tests to check the following for symmetry with respect to the axes and the origin. 2x-8^20=0 "^20=power of
Answer by stanbon(75887)

Use algebraic tests to check the following for symmetry with respect to the axes and the origin. 2x-8^20=0 "^20=power of
------------------------
f(x) = 2x-8^20
---
f(-x) = 2(-x)-8^20 = -2x-8^20
---
-f(-x) = -[-2x-8^20] = 2x+8^20
----
f(x) does not equal f(-x) so f is not symmetric to the y-axis
---
f(x) does not equal -f(-x) so f is not symmetric to the origin
-----------
Cheers,
Stan H.